OpenStudy (anonymous):

What is the process to simplifying 2√126 / √14 ?

zepdrix (zepdrix):

This is our problem? :) \[\Large \frac{2\sqrt{126}}{\sqrt{14}}\]

OpenStudy (anonymous):

Yes. Sorry I don't know how to type that into this on my iPod. But yes that is my problem.

zepdrix (zepdrix):

I guess this is one way we can rewrite our expression:\[\Large 2\sqrt{\frac{126}{14}}\]

zepdrix (zepdrix):

Any confusion on that step? :O It should be a piece of cake from there :) hehe First try it on your calculator is see if 14 divides evenly into 126.

OpenStudy (anonymous):

Yeah. It would equal 9. So I would end up with 2√9 ?

zepdrix (zepdrix):

Hmm good good. Looks like we can simplify a bit further though!! sqrt{9}=?

OpenStudy (anonymous):

3 of course :p so the whole thing simplified is 2√3, right.

OpenStudy (anonymous):

Or just 2 *3?

zepdrix (zepdrix):

2*3, yes :)

zepdrix (zepdrix):

which simplifies a bit further!!! lol

OpenStudy (anonymous):

Are you serious? (.0.) I thought I'd be done after that.

OpenStudy (anonymous):

Oh. I gotta multiply it right? x3

zepdrix (zepdrix):

That last step is a doozy, careful! 2*3=?